### v1

# Audio xLSTMs: Learning Self-supervised audio representations with xLSTMs

Abstract
While the transformer has emerged as the eminent neural architecture, several independent lines of research have emerged to address its limitations. Recurrent neural approaches have also observed a lot of renewed interest, including the extended long short-term memory (xLSTM) architecture, which reinvigorates the original LSTM architecture. However, while xLSTMs have shown competitive performance compared to the transformer, their viability for learning self-supervised general-purpose audio representations has not yet been evaluated. This work proposes Audio xLSTM (AxLSTM), an approach to learn audio representations from masked spectrogram patches in a self-supervised setting. Pretrained on the AudioSet dataset, the proposed AxLSTM models outperform comparable self-supervised audio spectrogram transformer (SSAST) baselines by up to 20% in relative performance across a set of ten diverse downstream tasks while having up to 45% fewer parameters.
Index Terms:
xLSTM, self-supervised learning, audio representation learning.
I Introduction
In lieu of their excellent generalisation capabilities and domain and data agnostic nature, transformers [1] and their successors have seen widespread adoption. Furthermore, transformers, together with masked predictive modelling, have emerged as a key driving force behind several prominent advancements in the realm of unsupervised and self-supervised representation learning for NLP [2], computer vision [3, 4] and audio and speech processing [5, 6, 7, 8, 9]. However, the race to find alternatives to scaled dot-product attention at the heart of transformers has heated up, with a lot of emphasis on finding sub-quadratic appromixations of the attention operation [10] as well as token mixing [11].
This search for alternatives has led us back to recurrent neural approaches. Before the advent of transformers, sequence modeling was predominantly done using recurrent neural networks (RNNs), such the long short-term memory (LSTM) [12]. Recurrent models offer several advantages over transformers: they scale linearly with respect to sequence length and they have lower runtime memory requirements since storing the entire key-value (KV) cache is not necessary. The search has led us to new approaches such as state space models (SSMs) [13, 14, 15, 16, 17], which are a family of sequence models that lie at the intersection of convolutional neural networks, RNNs and classical state spaces. Several variants of SSMs have since been proposed, showing competitive performance and scalability versus transformers in several domains, including long sequence modelling [16, 17], computer vision [18] as well as audio [19].
Compared to the transformer architecture, LSTMs suffer from several key drawbacks: (i) inability to revise storage decisions, (ii) compressing information into a scalar cell state, which impacts performance on rare input tokens, and (iii) memory mixing and the resulting lack of parallelizability. Recently, [20] proposed the extended long-short term memory (xLSTM) neural architecture, which revitalises the original LSTM architecture, leveraging the latest techniques and tricks learned from years of transformer and large language modelling research. xLSTM incorporates exponential gating, improved normalization and stabilization techniques while removing traditional memory mixing, resulting in two new fundamental building blocks: sLSTM and mLSTM. xLSTMs have demonstrated performance on-par with or better than transformers and Mamba for large language modelling [20] as well as a generic vision backbone trained on non-overlapping image patches [21], while also demonstrating superior sequence length extrapolation capabilities. However, the ability of xLSTMs to learn general-purpose audio representations in a self-supervised setting is yet to be thoroughly evaluated.
Figure 1: An overview of the proposed AxLSTM approach (left), and the constituent mLSTM blocks (right).
In this work, we propose Audio xLSTMs (AxLSTMs) for learning self-supervised general-purpose audio representations within a masked modelling framework from spectrograms patches. Pretrained on AudioSet [22], AxLSTMs consistently outperform comparable self-supervised audio spectrogram transformer (SSAST) [7] based baselines on ten varied downstream tasks, improving aggregate performance by over 20% while having over 45% fewer parameters. Code and pretrained models will be released publicly.
II Method
II-A Prerequisites: xLSTM and the mLSTM block
As previously discussed, xLSTM [20] proposes two new building blocks, sLSTM and mLSTM. In line with Vision-LSTM [21], the parallelizable mLSTM module is the fundamental building block of our approach, and thus is the focus of our discussion. The mLSTM block addresses the issues with the original LSTM cell by utilizing a matrix memory cell $C\in\mathbb{R}^{d\times d}$. mLSTM stores a key-value vector pair $k_{t},v_{t}\in\mathbb{R}^{d}$ at timestep $t$, and the relevant value vector is retrieved later using a query $q_{t+\tau}\in\mathbb{R}^{d}$. The forward pass for mLSTM is as follows:
${C_{t}}$ $=\ {f_{t}}\ {C_{t-1}}\ +\ {i_{t}}\ {v_{t}\ k_{t}^{\top}},$ (1)
${n_{t}}$ $=\ {f_{t}}\ {n_{t-1}}\ +\ {i_{t}}\ {k_{t}},$ (2)
$h_{t}$ $=\ {o_{t}}\ \odot\ {C_{t}}{q_{t}}\ /\ \max\left\{|{n_{t}^{\top}}{q_{t}}|,1\right\},$ (3)
$q_{t}$ $=\ W_{q}\ x_{t}\ +\ b_{q}$ (4)
$k_{t}$ $=\ \frac{1}{\sqrt{d}}W_{k}\ x_{t}\ +\ b_{k}$ (5)
$v_{t}$ $=\ W_{v}\ x_{t}\ +\ b_{v}$ (6)
${i_{t}}$ $=\ \exp{\left(\ w^{\top}_{i}\ x_{t}\ +\ b_{i}\right)}\ ,$ (7)
${f_{t}}$ $=\ \exp{\left(\ w^{\top}_{f}\ x_{t}\ +\ b_{f}\right)},$ (8)
$o_{t}$ $=\ \sigma\left(\ W_{o}\ x_{t}\ +\ b_{o}\right)\,$ (9)
where $C_{t}$ is the matrix memory cell, $n_{t}$ and $h_{t}$ represent the normalizer state and the hidden state, and $i_{t},f_{t}$ and $o_{t}$ represent the input, forget and the output gates, respectively. Weights $W_{q},W_{k}$ and $W_{v}$ are learnable projection matrices for vectors query $q$, key $k$ and value $v$, respectively. Since there is no memory mixing in mLSTMs, multiple memory cells and multiple heads are equivalent, and the forward pass can be parallelized. Further, mLSTM uses exponential gating to mitigate the inability of LSTMs to revise storage decisions.
II-B AxLSTM: Modelling masked patches with xLSTMs
Creating and masking patches: We compute non-overlapping patches of shape $t\times f$ from input spectrogram $\mathbf{x}\in\mathbb{R}^{\mathtt{t}\times\mathtt{f}}$, yielding $\mathbf{x}_{p}\in\mathbb{R}^{N\times(t\cdot f)}$ patches. After flattening these patches and projecting them linearly to a $\mathbb{R}^{N\times d_{m}}$ dimensional space, we encode positional information by adding fixed sinusoidal positional. A representative class token is then added to the beginning of the sequence, similar to [23, 7]. We then proceed to randomly mask 50% of the input patches using an unstructured masking strategy, and replace these masked patches with a learnable mask token. Thus, input to the encoder is:
$\mathbf{x^{\prime}}=[\mathtt{cls},\mathbf{x}_{p}^{1},\mathbf{x}_{p}^{2},\dots,\mathbf{x}_{p}^{N}]+E_{\mathtt{pos}}$ (10)
Encoding and Reconstruction: These partially masked patches are now fed to the encoder, yielding encoded representations $\mathbf{z}=\mathtt{enc}(\mathbf{x}^{\prime}),\mathbf{z}\in\mathbb{R}^{(N+1)\times d_{m}}$. The encoder is a stack of mLSTM blocks [20, 21] (Fig. 1), and each block expands the $d_{m}$ dimensional input by an expansion factor $E=2$ before projecting it back to $d_{m}$. This yields blocks that have fewer parameters than a corresponding transformer block. After encoding, a single hidden layer MLP is used to reconstruct patches from encoded representation $\mathbf{z}$. :
$\mathbf{y}^{\prime}$ $=\mathtt{Linear}_{(t\cdot f)}(\sigma(\mathtt{Linear}_{d_{m}}(\mathbf{z}))),$ (11)
where $\mathtt{Linear}_{d}$ is a parameterized linear projection to dimensions $d$, and $\sigma$ denotes the GELU non-linear activation function. The reconstructed output $y$ is obtained after removing the $\mathtt{cls}$ token from $\mathbf{y^{\prime}}$. The mean square error between the original input patches and the reconstructions is used for pretraining. Further details can be found in Section III-B.
III Experiments
TABLE I: Comparing AxLSTMs with popular self-supervised audio representations. We used pretrained models from cited papers to extract fixed feature vectors and conducted our own downstream experiments. Directly comparable SSAST and AxLSTM models are highlighted with similar colours. LS, AS, VP, LL stand for LibriSpeech, AudioSet, VoxPopuli and LibriLight datasets, respectively. SSAST [7] was trained on AS+LS. *includes decoder parameters
Model Data # Params BO CD ESC-50 LC Mri-S Mri-T NS-5h SC-5h F50K VL $s(m)$
Naive Baselines
HEAR-Naive [24] - - 52.6±2.4 30.9±0.8 5.8±0.2 33.5±1.1 38.0±1.3 36.4±1.9 18.6±4.4 8.5±0.4 7.1±0.2 11.2±0.5 5.2±0.8
Supervised
PaSST-Base [25] AS 86 M 94.9±0.5 61.0±0.3 94.8±0.3 60.1±0.2 96.5±0.1 87.6±0.6 23.3±0.9 66.6±1.4 64.2±0.1 25.5±0.8 74.7±0.4
SSL
W2V2-base [5] LS 94.4 M 74.0±1.0 46.4±0.3 31.1±0.4 51.2±0.2 77.3±0.2 55.1±0.3 7.4±0.8 90.8±0.3 18.1±0.1 35.5±0.8 44.1±0.2
W2V2-large [5] VP 315.4 M 93.1±0.7 66.9±0.4 60.1±0.5 62.4±0.3 93.9±0.1 77.4±0.2 42.0±1.0 87.6±0.5 34.2±0.1 53.6±1.0 75.1±0.4
WavLM-base [9] LS 94.4 M 89.4±0.7 56.3±0.2 46.6±0.4 63.2±0.3 95.1±0.1 83.4±0.2 37.3±0.8 57.2±0.8 29.9±0.1 22.6±0.6 61.7±0.2
WavLM-large [9] Mix 315.4 M 96.4±0.5 57.2±0.2 47.9±0.4 61.1±0.3 96.8±0.1 89.5±0.1 53.7±0.5 46.2±0.8 29.0±0.1 23.7±0.9 65.1±0.2
HuBERT-base [6] LS 94.4 M 92.1±0.6 70.8±0.2 57.8±0.6 56.5±0.3 94.4±0.1 84.9±0.3 19.4±0.7 93.2±0.1 32.3±0.1 61.8±0.6 74.5±0.2
HuBERT-large [6] LL 315.4 M 94.1±0.7 70.7±0.1 60.3±0.4 59.9±0.2 95.3±0.1 83.5±0.3 19.3±0.8 83.2±0.7 31.5±0.1 66.1±0.9 74.4±0.3
BEATs-iter3 [26] AS 90 M 94.0±0.8 67.3±0.2 83.7±0.3 68.0±0.2 94.7±0.1 95.8±0.1 69.4±0.8 85.2±0.3 53.6±0.2 38.5±1.0 87.6±0.3
AudioMAE [8] AS 86 M 93.7±0.6 68.2±0.2 60.6±0.4 42.2±0.2 89.2±0.2 86.6±0.2 64.5±0.8 28.6±1.5 37.9±0.1 29.7±1.0 64.3±0.3
MWMAE-Tiny [27] AS 12.6 M* 93.3±1.0 64.4±0.2 71.9±0.5 65.5±0.3 97.1±0.1 97.6±0.1 68.1±0.4 77.0±0.6 43.4±0.1 28.6±1.1 80.7±0.3
MWMAE-Base [27] AS 92.5 M* 96.0±0.5 73.1±0.3 81.2±0.4 68.8±0.2 97.4±0.1 97.9±0.1 69.3±0.6 90.9±0.2 51.2±0.2 44.2±0.9 91.3±0.2
SSAM-Tiny [19] AS 4.8 M 93.7±0.8 61.8±0.3 70.6±0.2 59.2±0.4 97.1±0.1 94.9±0.1 62.0±0.7 74.8±0.4 41.3±0.2 27.8±1.0 76.3±0.2
SSAM-Small [19] AS 17.9 M 94.0±0.7 67.5±0.2 78.7±0.6 60.5±0.3 97.5±0.1 96.7±0.1 66.3±0.8 83.7±0.3 48.5±0.1 39.6±0.7 84.4±0.3
SSAM-Base [19] AS 69.3 M 93.2±1.1 70.3±0.2 81.0±0.3 63.5±0.2 97.7±0.1 96.9±0.1 70.5±0.5 87.9±0.3 52.2±0.1 50.4±0.7 89.7±0.3
SSAST Based
SSAST [7] Mix 89 M 93.4±0.9 56.5±0.2 68.4±0.4 60.7±0.3 96.7±0.1 96.3±0.1 66.8±0.7 53.5±1.3 38.2±0.1 28.5±0.9 73.1±0.2
SSAST-Tiny AS 5.4 M 90.4±0.7 46.9±0.2 42.4±0.6 42.7±0.2 95.7±0.1 94.3±0.1 61.2±0.5 50.6±1.6 24.6±0.1 13.8±1.0 56.0±0.2
SSAST-Small AS 21.5 M 93.2±0.5 51.6±0.2 50.1±0.6 50.0±0.3 96.2±0.1 95.0±0.1 63.8±0.4 58.3±1.2 31.6±0.1 15.6±0.7 63.4±0.3
SSAST-Base AS 85.7 M 93.1±0.7 56.0±0.4 59.6±0.7 52.9±0.3 96.6±0.1 96.2±0.2 64.6±0.8 66.1±1.0 37.5±0.1 19.2±0.9 69.2±0.3
Proposed
AxLSTM-Tiny AS 2.9 M 93.1±0.7 53.7±0.4 58.3±0.3 54.2±0.4 96.4±0.1 94.2±0.1 61.6±0.7 62.5±1.1 35.1±0.1 20.4±1.0 67.0±0.3
AxLSTM-Small AS 11.2 M 94.2±0.4 63.4±0.2 72.2±0.5 61.7±0.4 97.2±0.0 96.4±0.1 71.7±0.4 76.0±0.9 43.6±0.1 31.3±0.5 79.8±0.2
AxLSTM-Base AS 44.0 M 93.9±0.6 66.1±0.2 75.9±0.5 63.0±0.3 97.0±0.1 96.9±0.1 72.1±0.6 80.4±0.3 47.1±0.1 35.8±0.8 83.1±0.2
III-A Datasets
Pretraining: For pretraining, we use the AudioSet dataset (AS) [22] which has roughly 2 million 10-second weakly labelled YouTube clips and over 5000 hours of audio data spanning 527 classes.
Downstream Evaluation: Following recent works [27, 19], we use a subset of the tasks proposed as a part of the HEAR benchmark [24] for downstream evaluation, comprising of the following ten diverse audio tasks: Beijing Opera, Crema-D, ESC-50, LibriCount, Mridangam Stroke and Tonic, NSynth Pitch 5h, Speech Commands 5h, FSD50K and VoxLingua107.
III-B Implementation details
Spectrogram features: We extract log-scaled mel spectrogram features with a window size of $25$ ms, a hop size of $10$ ms and $F=80$ mel-spaced frequency bins. A sampling rate of $16000$ Hz is used for all audio clips.
Pretraining: All the proposed models are pretrained on randomly cropped 2-second audio clips, which corresponds to a [200, 80]-dim (time and frequency, resp.) input. Our default configuration consists of $l=12$ number of stacked mLSTM blocks with a model feature dimension $d_{m}=192$, same as those of ViT-Tiny [23] encoder, but we also evaluate Small ($d_{m}=$384, $l$=12) and Base ($d_{m}$=768, $l$=12) encoder configurations. The default expansion factor $E_{f}=2$ as specified in [20, 21] is used in the underlying mLSTM layers. By default, a patch embedding layer that computes ($4\times 16$) shaped non-overlapping patches is used in all our experiments. All models are trained for 100 epochs with a batch size of 1024 and a weight decay of 0.05. AdamW optimizer with a linear warmup for 10 epochs followed by a cosine learning rate decay schedule is used. No data augmentations were used. We follow hyperparameter recommendations from previous work [19, 7].
Downstream evaluation: Following the HEAR protocol, we extract fixed sized feature vectors independent of the input audio duration by taking the mean over time across 2-second audio chunks. We then train a single hidden layer MLP classifier with 1024 neurons for each task, using the official hear-eval-kit accompanying the HEAR benchmark. All experiments are repeated with 10 different random seeds.
Aggregated Performance Metric: We use the aggregated normalized score as proposed by [27] to compare evaluated approaches across the proposed list of downstream tasks. For a model $m$, overall score $s(m)\in[0.,100.]$ is given as: $s(m)=\frac{1}{|T|}\sum_{t\in T}\frac{x_{t}(m)-min_{t}}{max_{t}-min_{t}}*100$, where $x_{t}(m)$ denotes performance of the model $m$ on task $t$, and $min_{t}$ and $max_{t}$ represent the worst and the best performance across all models on the task, thus taking into account the relative performance amongst all evaluated representations.
IV Results
IV-A Comparison with existing works
Table I shows how the proposed AxLSTM models fare against recent audio representations. Directly comparable SSAST and AxLSTM feature representations are colour coded and have identical feature embedding sizes, however, feature vector sizes extracted using other referred methods can vary. While sub-optimal, it is infeasible to retrain all pretrained representations to have the same embedding sizes, and our evaluation protocol is in line with recent frameworks for evaluating self-supervised audio representations [24, 28]. SSAST [7] represents the official SSAST released model, which was pretrained on AudioSet and LibriSpeech datasets with the masked prediction + reconstruction multitask objective. We can see that AxLSTM models consistently outperform their transformer based SSAST counterparts by a considerable margin, while having over 45% fewer parameters, with AxLSTM-Base configuration yielding a 30% relative improvement in aggregate performance (83.1±0.2 v/s 69.2±0.3). While Mamba based SSAM [19] as well as Masked Autoencoder based approaches [27] perform better, it is worth noting that the proposed AxLSTM models have about 35% and 50% fewer parameters, respectively. Overall, we can conclude that AxLSTMs performs very favourably compared to popular audio representations.
IV-B Ablations
Expansion Factor: We investigate how AxLSTM models scale with different expansion factors ($E_{f}$). Table II shows that while using $E_{f}=3$ improves performance, $E_{f}=4$ leads to worse performance, suggesting overfitting.
TABLE II: Performance impact of expansion factor ($E_{F}$)
Model $E_{f}$ Params $s(m)$
AxLSTM-Tiny 2 2.9 M 67.0±0.3
AxLSTM-Tiny 3 4.3 M 68.9±0.4
AxLSTM-Tiny 4 5.8 M 63.4±0.2
Patch Size: To investigate how AxLSTM models perform compared to SSAST models with changing number of input patches, we pretrain AxLSTM-Tiny models with 3 patch sizes: $(4,8)$, $(4,16)$, and $(8,16)$. AxLSTM scales better with increasing number of patches, with smaller patch sizes leading to better performance (Table III). We expect expansion factor $E_{f}=3$ paired with a patch size of $(4,8)$ to further improve performance.
TABLE III: Patch size ablations with the Tiny configuration
Model Patch Size # Patches $s(m)$
SSAST-Tiny $(8,16)$ 125 47.3±0.3
AxLSTM-Tiny $(8,16)$ 125 55.3±0.3
SSAST-Tiny $(4,16)$ 250 56.0±0.2
AxLSTM-Tiny $(4,16)$ 250 67.0±0.3
SSAST-Tiny $(4,8)$ 500 54.1±0.4
AxLSTM-Tiny $(4,8)$ 500 68.3±0.3
Other design choices: There are several design choices that impact the modeling performance of AxLSTMs. In Table IV, we evaluate the impact of two of the most important of these factors using the AxLSTM-Tiny configuration. We observe that the absence of exponential gating for the input and forget gates in mLSTM ends up improving performance. This can be attributed to the fact that for the given input sequence length (250 tokens), revising storage decisions is not a bottleneck. Further, we observe that removing sequence flipping in even numbered blocks [21] significantly boosts performance, highlighting that bidirectional processing of patches is not as important in the given framework as it is for vision, which is in line with observations made in previous works [19].
TABLE IV: Evaluating other design choices for AxLSTM-Tiny
$s(m)$
default 67.0±0.3
default, no exponential gating 68.2±0.2
default, no sequence flipping 70.2±0.3
V Conclusion
This work presents extended long short-term memory models for audio (AxLSTMs), an approach for learning self-supervised general-purpose audio representations from masked audio spectrograms which extends the recently proposed xLSTM neural architecture. Using the AudioSet corpus for pretraining, we evaluate AxLSTMs against directly comparable Self-supervised Audio Spectrogram transformer (SSAST) baselines on ten varied downstream audio recognition tasks. Our empirical analysis shows that AxLSTMs offer up to 20% improvement in relative performance over their transformer based counterparts while having up to 45% fewer parameters, thus establishing the viability of xLSTM approaches for learning self-supervised audio representations.
Acknowledgment
We would like to thank Prof. Lars Kai Hansen, Technical University of Denmark, for valuable feedback. We are also grateful to the TPU Research Cloud Program, a Google Research Initiative, for providing TPU devices used in this project, as well as the Danish e-Infrastructure Consortium (DeiC) for enabling access to the LUMI supercomputer.